O'Brien Ltd.'s outstanding bonds have a $1,000 par value, and they mature in 25 years. Their nominal yield to maturity is 9.25%, they pay interest semiannually, and they sell at $975. What is the bond's nominal coupon interest rate?

Please show all work, without excel
Answer #1

Number of periods = 25 * 2 = 50

Rate = 9.25% / 2 = 4.625%

Price = Coupon * [ 1 - 1 / ( 1 + r)n] / r + FV / ( 1 + r)n

975 = Coupon * [ 1 - 1 / ( 1 + 0.04625)50] / 0.04625 + 1000 / ( 1 + 0.04625)50

975 = Coupon * 19.366784 + 104.2861

870.7139 = Coupon * 19.366784

Coupon = 45

Annual coupon = 45 * 2 = 90

Coupon interest rate = ( 90 / 1000) * 100

Coupon interest rate = 0.09 or 9%